OpenStudy (anonymous):

solve 5/4x + 1/8 = 9/8 +x the solution is

OpenStudy (anonymous):

1/4x=1 x=4

OpenStudy (anonymous):

I left out an x on the problem so i resend 5/4x + 1/8x = 9/8 + x the solution is x = ?

OpenStudy (anonymous):

\[\frac{5 x}{4}+\frac{x}{8}\text{=}\frac{9}{8}+x \]x=3